TIME-RESOLVED FTIR EMISSION SPECTROSCOPY OF THE ν_{1} CH STRETCH MODE OF THE KETENYL (HCCO) RADICAL

\underline{M. WILHELM}, W. MCNAVAGE, Department of Chemistry, University of Pennsylvania, Philadelphia, PA 19104; H. L. DAI, Department of Chemistry, Temple University, Philadelphia, PA 19122.

The ν_{1} CH stretch, a previously uncharacterized mode of the ketenyl (HCCO) radical, was detected at 3230cm^{−1} through time-resolved Fourier transform infrared emission spectroscopy. Ro-vibrationally excited ketenyl and ethyl (CH_{2}CH_{3}) radicals were generated, with near unit quantum efficiency, via the 193 nm photodissociation of ethyl ethynyl ether ^{a}.

HCC-O- CH_{2}CH_{3} + hν (193 nm) → HCCO + CH_{2}CH_{3}

IR emission from the photoproducts was detected with both temporal and frequency resolution. Spectral assignments were made based upon comparison with theoretical calculations as well as 2D correlation analysis^{b}.

aM. J. Krish, J. L. Millerm, L. J. Butler, H. Su, R. Bersohn, and J. Shu, J. Chem. Phys. 119, 176(2003).

bW. McNavage, and H. L. Dai, J. Chem. Phys. 123, 184104(2005).